OpenStudy (anonymous):

complete the square add number to both sides x^2-6x=-8

OpenStudy (anonymous):

do you want to find x here? :)

OpenStudy (nikvist):

\[x^2-6x=-8\quad\Rightarrow\quad x^2-6x+9=-8+9\quad\Rightarrow\quad (x-3)^2=1\]